LTE stock sells for $27.69 per share and there are 280,000 shares outstanding. The company plans a 4-for-3 reverse stock split. Assuming no market imperfections or tax effects, what will the stock price be after the split?

.
Answer #1

Stock will be reverse Split i.e. every 4 shares will be converted to 3 share.

Value of Shares = No of Shares * Value per Share

= 280,000 * 27.69

= $ 7,753,200

So new no of shares will be 280,000 * 3 /4 = 210,000Shares after reverse Split

Value after reverse split remians same So, Share Price will be

$ 7,753,200 = New No of shares * Share Price

$ 7,753,200 = 210,000 * Share Price

Share Price = $ 7,753,200/ 210,000 = $ 36.92

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you